OpenStudy (anonymous):

solve: 4-(1/t)-(3/t^2)=0 t= ?? simplify as integer or fraction, using radicals as needed. express complex number in terms of i.

jimthompson5910 (jim_thompson5910):

Start by multiplying EVERY term by the LCD t^2 to get 4t^2-t-3=0 Now let's factor and solve (t-1)(4t+3)=0 t-1=0 or 4t+3=0 t=1 or t=-3/4